Thursday, October 12, 2017

- Aluminum plug was made and placed in sight glass hole of oil trap.

- Aluminum plug was push out of oil trap by built up pressure.
- Clamp was used to secure Al plug in place while emptying H2 lines.

- Ran air through oil trap and scrubber.  Measured a 40psi pressure build up that would slowly decrease when air supply was stopped.
- Removed scrubber, emptied charcoal and measured ~3psi pressure build up while pushing air through the empty scrubber.
- Cleaned inside of scrubber with IPA and let dry overnight.

- Filled scrubber with new charcoal and measured ~10psi pressure build up when pushing air through it.
- Reconnected scrubber and oil trap still have clamp on oil trap plug, since we could not remove it.  However, oil trap was pressurized to ~70psi and the Al plug did not move. 

Tuesday, October 3, 2017

- Modified oil trap so both large and small filters can be used in it.
- Changed oil trap filter, using a larger one than normal.

Friday, September 29, 2017

- Small leak detected at the sight glass on the oil trap.  Broke spanner while trying to reposition it.  Fixed sight glass in place with wire and super glue.

- Still picking up a small leak from the sight glass of the oil trap.  Covered it with vacuum epoxy.

Thursday, September 28, 2017

- Changed oil purifier filters and pump oil.

- Drained oil from pump and opened and cleaned oil tray and sight glass in the pump.  Refilled with new oil and ran vacuum pump and oil purifier the remainder of the day,  TGM was reading less than 5 ppb till around 8pm at which point it dropped to 0 ppb.  Pump and purifier were shut off at 9pm.

-Tried a growth and alarm went off at "fill lines" step.  Let pump cabinet air out then used hand sensor to test for leaks while running H2.  Hand sensor picked up on a leak near the pump cap, which also triggered the hand sensor without any H2 flowing through the system.  Left TGM tape around multiple points on the pump over night and the tape around the pump cap became discolored.

- Changed the O-ring on the pump cap as well as wrapped the cap threads with Teflon to create a better seal between the cap and pump.  Hand sensor didn't pick anything up so H2 and AsH3 were flushed through the system for 10min with no reading on the TGM.

- Tape was also changed on the TGM.

Friday, September 22, 2017


TGM alarm maxed out first in pump enclosure then in both pump and reactor enclosures, with minimum readings at the scrubber outlet and system enclosure.

Found a crack pipe clamp on the vacuum line going to the pump.  Replaced this connecting pipe clamp.

Replaced the pump oil.

Wednesday, June 21, 2017

AsH3 cylinder was replaced. 
     Lot # 1987306991Q7  SV020731
     Net Weight: 1.98 lbs or 898.11 g

Old cylinder is being stored in back of MOCVD enclosure.

Wednesday, May 10, 2017


Charcoal in scrubber was changed.  This was done to help prevent readings at the scrubber outlet when using PH3.


A finer grade of charcoal was added to the charcoal in the scrubber.

TGM pump stopped working during the night.  A fastener screw was sheared off and I was unable to adjust the flow at any of the points.  All points were reading low flow.  The pump was replaced with a new one, which is working fine. - NS